### v1

# Equivalent Systems for Differential Equations with Polynomially Distributed Delay

Abstract
We consider delay differential equations with a polynomially distributed delay. We derive an equivalent system of delay differential equations, which includes just two discrete delays. The stability of the equivalent system and its stationary solutions are investigated. Alternatively, a Gaussian quadrature generates a discretisation of the integral, which describes the distributed delay in the original delay differential equation. This technique yields an approximate differential equation with multiple discrete delays. We present results of numerical computations, where initial value problems of the differential equations are solved. Therein, the two approaches are compared.
Keywords: delay differential equation, distributed delay, initial value problem, stationary solutions, stability, Gaussian quadrature.
1 Introduction
Mathematical modelling by delay differential equations (DDEs) is used in many applications like epidemiology, population dynamics, traffic flow, and others, see [5, 13]. A DDE includes discrete delays and/or distributed delays. Numerical methods were developed for the solution of initial value problems (IVPs) of DDEs with discrete delays, see [1, 8]. In the case of a distributed DDE with a gamma distribution or a similar distribution, it is well known that an equivalent system of ordinary differential equations (ODEs) can be arranged, see [4, 6, 9]. Thus numerical methods for IVPs of ODEs are applicable, see [3].
We investigate a DDE including a distributed delay, where the weight function of the distribution is a polynomial. Such a polynomial is present in a uniform distribution and a beta distribution, for example. We construct an equivalent system of DDEs with just two discrete delays. We examine the stability of this system, where we focus on stationary solutions. Furthermore, a scaling of the system is formulated, which keeps the magnitude of its solutions small, such that error amplification and bad condition are reduced.
Alternatively, the integral defining a distributed delay can be discretised by a quadrature method, see [11]. This discretisation yields a DDE with multiple discrete delays, whose solution represents an approximation of the solution for the distributed DDE. We employ a Gaussian quadrature associated to the weight function, see [14], in this technique.
Finally, we demonstrate results of numerical computations using a nonlinear system of DDEs with distributed delay as test example. Therein, the two approaches, i.e., equivalent DDE system and approximate DDE system, are compared. The IVPs are solved by a numerical method for DDEs with multiple discrete delays, see [12].
2 Problem definition
We consider an autonomous DDE of the form
$y^{\prime}(t)=f\Big{(}y(t),\int_{I}y(t-\tau)g(\tau)\;\mathrm{d}\tau\Big{)}$ (1)
with a continuous function $f:\mathbbm{R}\times\mathbbm{R}\rightarrow\mathbbm{R}$. The interval $I$ is either the unbounded domain $I=[0,\infty)$ or a compact domain $I=[a,b]$, $0\leq a<b<\infty$. The bounds $a=\tau_{\min}$ and $b=\tau_{\max}$ represent a minimum delay and a maximum delay, respectively. The weight function $g:I\rightarrow\mathbbm{R}$ has the properties of a probability density function:
  • i)
    $g$ is measurable
  • ii)
    $g(\tau)\geq 0$  for all $\tau\in I$,
  • iii)
    $\int_{I}g(\tau)\;\mathrm{d}\tau=1$  (normalisation).
The condition (iii) is imposed without loss of generality, since a normalising constant of the weight function $g$ could be included in the right-hand side $f$. It is just required that the integral of $g$ is finite.
Remark 1
We study the case of a single DDE (1). Nevertheless, generalisations to systems of DDEs with distributed delay are straightforward. Furthermore, the procedure can be applied to non-autonomous DDEs as well.
Solutions of IVPs (with $I=[a,b]$) are characterised as follows.
Definition 1
Let $\phi:[-b,0]\rightarrow\mathbbm{R}$ be a continuous initial function. A continuous function $y:[-b,\infty)\rightarrow\mathbbm{R}$ is a solution of the IVP of the DDE (1), if it holds that
  • i)
    $y$ is continuously differentiable on $(0,\infty)$ and $\lim_{t\rightarrow 0+}y^{\prime}$ exists,
  • ii)
    $y$ satisfies the DDE for $t>0$,
  • iii)
    $y(t)=\phi(t)$ for $t\in[-b,0]$.
In the case of a gamma distribution, the probability density function reads as
$g(\tau)=C\,x^{p}\mathrm{e}^{-\tau/\theta}\qquad\mbox{for}\;\;\tau\geq 0$
with parameters $\theta,p>0$ and a normalising constant $C>0$. Let $p\geq 1$ be an integer. A linear chain technique yields an equivalent system of ODEs associated to the DDE (1), where the solution $y$ is one component in a solution of the ODEs, see [4, 9]. Consequently, numerical methods for ODEs can be used, see [3].
We study the case of a polynomial $g$ on a compact interval $[a,b]$. Thus the weight function exhibits the general form
$g(\tau)=\sum_{i=0}^{n}\alpha_{i}\tau^{i}$ (2)
with $\alpha_{n}\neq 0$. An important class of probability distributions with polynomial density is the beta distribution. The probability density function
$g(\tau)=C\,(\tau-a)^{p}\,(b-\tau)^{q}\qquad\mbox{for}\;\;\tau\in[a,b]$ (3)
depends on parameters $p,q\geq 0$, whereas $C>0$ is a normalising constant. The uniform distribution ($g$ constant) represents the special case of $p=q=0$.
3 Equivalent systems
We investigate the existence of equivalent systems for the DDE (1) in the case of a polynomially distributed delay with (2).
3.1 Preliminaries
The construction of an equivalent system is based on the following property.
Theorem 1
Let $y:[-b,\infty)\rightarrow\mathbbm{R}$ be continuous. We define
$x_{i}(t)=\int_{a}^{b}y(t-\tau)\tau^{i}\;\mathrm{d}\tau$ (4)
for $i\geq 0$ and $t\geq 0$. It follows that $x$ is continuously differentiable and
$x_{i}^{\prime}(t)=y(t-a)a^{i}-y(t-b)b^{i}+ix_{i-1}(t)$
for $i\geq 0$ and $t>0$, without loss of generality let $x_{-1}\equiv 0$.
Proof:
Using the substitution $s=t-\tau$, the integral in (4) changes into the form
$x_{i}(t)=\int_{t-b}^{t-a}y(s)\,(t-s)^{i}\;\mathrm{d}s$
for $i\geq 0$. In the case of $i=0$, the fundamental theorem of calculus yields
$x_{i}^{\prime}(t)=y(t-a)-y(t-b)=y(t-a)a^{0}-y(t-b)b^{0}.$
The binomial formula implies for $i\geq 1$
$x_{i}(t)$ $=\int_{t-b}^{t-a}y(s)\,\sum_{k=0}^{i}\binom{i}{k}t^{k}(-s)^{i-k}\;\mathrm{d}s\;=\;\sum_{k=0}^{i}\binom{i}{k}t^{k}\int_{t-b}^{t-a}y(s)\,(-s)^{i-k}\;\mathrm{d}s$
$=\int_{t-b}^{t-a}y(s)\,(-s)^{i}\;\mathrm{d}s+\sum_{k=1}^{i}\binom{i}{k}t^{k}\int_{t-b}^{t-a}y(s)\,(-s)^{i-k}\;\mathrm{d}s.$
Now the integrands are continuous and independent of $t$. Thus differentiation with respect to $t$ can be done using the fundamental theorem of calculus. We obtain with the product rule of differentiation
$x_{i}^{\prime}(t)$ $=y(t-a)(a-t)^{i}-y(t-b)(b-t)^{i}$
$\qquad+\sum_{k=1}^{i}\binom{i}{k}kt^{k-1}\int_{t-b}^{t-a}y(s)\,(-s)^{i-k}\;\mathrm{d}s$
$\qquad+\sum_{k=1}^{i}\binom{i}{k}t^{k}\big{(}y(t-a)(a-t)^{i-k}-y(t-b)(b-t)^{i-k}\big{)}$
$=y(t-a)a^{i}-y(t-b)b^{i}+\sum_{k=1}^{i}\binom{i}{k}kt^{k-1}\int_{t-b}^{t-a}y(s)\,(-s)^{i-k}\;\mathrm{d}s.$
We examine the sum further
$\sum_{k=1}^{i}\binom{i}{k}kt^{k-1}\int_{t-b}^{t-a}y(s)\,(-s)^{i-k}\;\mathrm{d}s$
$=$ $\;i\sum_{k=1}^{i}\frac{(i-1)!}{(k-1)!(i-k)!}\,t^{k-1}\int_{t-b}^{t-a}y(s)\,(-s)^{i-k}\;\mathrm{d}s$
$=$ $\;i\sum_{k=0}^{i-1}\frac{(i-1)!}{k!(i-1-k)!}\,t^{k}\int_{t-b}^{t-a}y(s)\,(-s)^{i-1-k}\;\mathrm{d}s$
$=$ $\;i\int_{t-b}^{t-a}y(s)\sum_{k=0}^{i-1}\binom{i-1}{k}t^{k}(-s)^{i-1-k}\;\mathrm{d}s\,=\,i\int_{t-b}^{t-a}y(s)\,(t-s)^{i-1}\;\mathrm{d}s\,=\,ix_{i-1}(t).$
Hence the formula of Theorem 1 is shown. $\Box$
Remark 2
If the function $y$ is continuously differentiable everywhere, then the proof of Theorem 1 is straightforward by interchanging differentiation and integration. It is surprising that just the continuity of $y$ is sufficient for the differentiability of the functions (4).
3.2 Derivation of systems
In the case of a polynomial weight function (2), the integral of the DDE (1) becomes
$\int_{a}^{b}y(t-\tau)\,g(\tau)\;\mathrm{d}\tau=\sum_{i=0}^{n}\alpha_{i}\int_{a}^{b}y(t-\tau)\,\tau^{i}\;\mathrm{d}\tau=\sum_{i=0}^{n}\alpha_{i}x_{i}(t)$ (5)
including the functions (4). Using Theorem 1, we arrange the following system of DDEs.
Definition 2
Let (1) be a DDE including a polynomially distributed delay, see (2). An equivalent system of DDEs with two discrete delays is
$y^{\prime}(t)$ $=f\Big{(}y(t),\sum_{i=0}^{n}\alpha_{i}x_{i}(t)\Big{)}$ (6a)
$x_{0}^{\prime}(t)$ $=y(t-a)-y(t-b)$ (6b)
$x_{i}^{\prime}(t)$ $=y(t-a)a^{i}-y(t-b)b^{i}+ix_{i-1}(t)\quad\mbox{for}\;\;i=1,\ldots,n.$ (6c)
In an IVP of the distributed DDE (1), the initial conditions $y(t)=\phi(t)$ for $t\in[-b,0]$ are predetermined. Concerning the variables $x_{i}$, initial conditions are required only at $t=0$. Due to (4), it follows that
$x_{i}(0)=\int_{a}^{b}\phi(-\tau)\tau^{i}\;\mathrm{d}\tau$ (7)
for $i=0,1,\ldots,n$. Thus the initial values are computable. Often constant initial values $y(t)=y_{0}$ for $t\leq 0$ are employed. In this case, we obtain
$x_{i}(0)=y_{0}\,\frac{b^{i+1}-a^{i+1}}{i+1}$ (8)
for $i=0,1,\ldots,n$. Now numerical methods for IVPs of DDEs with a finite number of discrete delays can be used, see [1].
The following theorem shows the equivalence of the differential equations.
Theorem 2
Let $y$ be a solution of an IVP of the DDE (1). It follows that $y$ and the additional functions (4) satisfy the system (6). Vice versa, let $y,x_{0},\ldots,x_{n}$ be a solution of an IVP of the system (6) with initial values $\phi$ for $y$ and (7) for $x_{i}$. Consequently, the part $y$ solves the DDE (1) with initial values $\phi$.
Proof:
We consider a solution $y$ of an IVP of the DDE (1). Thus $y$ satisfies the first equation in (6) due to the definition of the additional functions (4). Theorem 1 guarantees that the functions (4) satisfy Eqns. (6b) and (6c).
Alternatively, we examine a solution $y,x_{0},\ldots,x_{n}$ of an IVP of the system (6). We consider Eqns. (6b) and (6c) for the unknowns $x=(x_{0},x_{1},\ldots,x_{n})^{\top}$ and the fixed function $y$. Hence (6b), (6c) represents a system of ODEs for $x$, where the right-hand side is continuous and continuously differentiable with respect to $x$. Thus solutions of IVPs are unique. Using the initial values (7), the solution of (6b), (6c) are the functions (4) due to Theorem 1. The relation (5) shows that Eq. (6a) is equivalent to the DDE (1). $\square$
Remark 3
In Theorem 2, it is not required that solutions of IVPs of the DDE (1) are unique.
3.3 Stability
We investigate the stability of the equivalent system and its stationary solutions. In the system (6), the parts (6b), (6c) exhibit the form
$x^{\prime}(t)=Ax(t)+u(y(t-a),y(t-b))$ (9)
with the variables $x=(x_{0},x_{1},\ldots,x_{n})^{\top}$, a matrix
$A=\begin{pmatrix}0&&&&\\ 1&0&&&\\ &2&\ddots&&\\ &&\ddots&0&\\ &&&n&0\\ \end{pmatrix},$ (10)
and a linear function $u$. The rank of the matrix (10) is $n$. The matrix (10) has the sole eigenvalue $\lambda=0$ with algebraic multiplicity $n+1$. There is just a single eigenvector $e=(0,\ldots,0,1)^{\top}$. Thus the geometric multiplicity of the eigenvalue is one. It follows that the linear ODE
$\hat{x}^{\prime}(t)=A\hat{x}(t)$ (11)
is Lyapunov-stable for $n=0$ and unstable for $n\geq 1$, see [2, p. 376]. Hence the linear term $Ax$ induces an instability in the system (6). However, error terms just grow polynomially and not exponentially in time, since there is no eigenvalue with a positive real part. In view of this instability, IVPs of a system (6) should be solved with high accuracy requirements to reduce the growth of errors.
Stationary solutions of a DDE are of the form $y(t)=y^{*}$ for all $t$ with a constant $y^{*}$. A constant $y^{*}$ represents a stationary solution of (1), if and only if the algebraic equation
$f(y^{*},y^{*})=0$
is satisfied. The associated stationary solutions of the additional functions (4) read as
$x_{i}^{*}=\int_{a}^{b}y^{*}\tau^{i}\;\mathrm{d}\tau=y^{*}\,\frac{b^{i+1}-a^{i+1}}{i+1}$ (12)
for $i=0,1,\ldots,n$, like in the calculation of (8).
If $y^{*}$ is an asymptotically stable stationary solution, then we obtain the convergence
$\lim_{t\rightarrow 0}x_{i}(t)=\lim_{t\rightarrow\infty}\int_{a}^{b}y(t-\tau)\tau^{i}\;\mathrm{d}\tau=\int_{a}^{b}\lim_{t\rightarrow\infty}y(t-\tau)\tau^{i}\;\mathrm{d}\tau=x_{i}^{*}$
for $i\geq 0$ provided that the initial values of $y$ are sufficiently close to $y^{*}$ and the initial values (7) are used for $x_{i}$. An interesting question is if this stationary solution of (6) is asymptotically stable or Lyapunov-stable.
Theorem 3
We suppose the existence and uniqueness of solutions of IVPs for all $t\geq 0$. Let $y^{*}$ be a Lyapunov-stable stationary solution of a DDE (1). Consequently, $(y^{*},x_{0}^{*},\ldots,x_{n}^{*})^{\top}$ with $x_{i}^{*}$ from (12) is a stationary solution of the equivalent system (6), which is unstable for $n\geq 2$.
Proof:
We suppose $y^{*}=0$ without loss of generality. It follows that $x_{i}^{*}=0$ for all $i$. The Lyapunov-stability implies that for an $\varepsilon>0$, there is a $\delta\in(0,\varepsilon]$ such that the solution of an IVP is bounded by $|y(t)|<\varepsilon$ for all $t>0$, if the initial function satisfies $|\phi(t)|<\delta$ for all $t\leq 0$.
Let $z=(y,x_{0},\ldots,x_{n})^{\top}$ be a solution of the equivalent system (6) with initial functions $z_{0}=(\phi,\psi_{0},\ldots,\psi_{n})^{\top}$. Let $\tilde{\delta}\in(0,\delta]$. If $z_{0}$ is in a $\tilde{\delta}$-neighbourhood of zero with respect to the maximum norm, then the part $\phi$ is in a $\delta$-neighbourhood of zero. In the right-hand side of (6b), (6c), we estimate the terms
$\left|y(t-a)a^{i}-y(t-b)b^{i}\right|\leq(a^{n}+b^{n})\sup_{s\geq-b}|y(s)|\leq(a^{n}+b^{n})\varepsilon$
for $i=0,1,\ldots,n$ and $t\geq 0$. Thus it holds that $\|u\|_{\infty}\leq(a^{n}+b^{n})\varepsilon$ for the function $u$ in (9). We consider the two systems (9) and (11) with same initial conditions. The matrix (10) is nilpotent with $A^{n}\neq 0$ and $A^{n+1}=0$. Thus its matrix exponential reads as
$\mathrm{e}^{tA}=\sum_{j=0}^{n}\frac{t^{j}}{j!}\,A^{j}.$
All terms are non-negative for $t\geq 0$ in this sum. We consider solutions of the system (11) in the case of $n\geq 2$. In each $\tilde{\delta}$-neighbourhood of zero, there are initial conditions such that $\|\hat{x}(t)\|_{\infty}\geq Kt^{n}\geq Kt^{2}$ for all $t\geq 1$, where the constant $K>0$ depends on the initial values. In contrast, it holds that
$\|x(t)-\hat{x}(t)\|_{\infty}\leq t\sup_{s\in[0,t]}\|u(y(s-a),y(s-b))\|_{\infty}\leq(a^{n}+b^{n})\varepsilon\,t$
for any $t\geq 0$. We obtain a chain of inequalities
$Kt^{2}\leq\|\hat{x}(t)\|_{\infty}\leq\|x(t)-\hat{x}(t)\|_{\infty}+\|x(t)\|_{\infty}\leq(a^{n}+b^{n})\varepsilon\,t+\|x(t)\|_{\infty}.$
It follows that
$(Kt^{2}-(a^{n}+b^{n})\varepsilon t)\leq\|x(t)\|_{\infty}$
for all $t\geq 1$. Now $\|x(t)\|_{\infty}$ is not bounded for increasing $t$. Hence the stationary solution zero is not Lyapunov-stable. $\square$
Remark 4
The above proof of Theorem 3 does not work in the case of $n=1$, because a linear growth of a solution $\hat{x}$ of (11) could be compensated in a solution $x$ of (9) by the term $u$. However, it is very unlikely that the term $u$ balances the term $Ax$ for all positive times. Thus we expect that the stationary solution is unstable in the case $n=1$ as well.
3.4 Scaling
Concerning the equivalent system (6), the additional functions (4) include powers $\tau^{i}$. It follows that the functions (4) exhibit huge magnitudes for large delays. This effect may be harmful due to the amplification of errors. However, the disadvantageous effect can be avoided by a scaling in the original DDE (1).
We introduce the new time variable $\tilde{t}=t/b$. Since it holds that $b=\tau_{\max}$, all involved delays are smaller or equal one in the new variable. In particular, it holds that $10<\frac{a}{b}<$ 1. Now the DDE (1) is transformed into
$\tilde{y}^{\prime}(\tilde{t})=b\,f\Big{(}\tilde{y}(\tilde{t}),\int_{a/b}^{1}\tilde{y}(\tilde{t}-\tau)\tilde{g}(\tau)\;\mathrm{d}\tau\Big{)}.$ (13)
The new weight function $\tilde{g}$ is just the linear transformation of $g$ from the interval $[a,b]$ to the interval $[\frac{a}{b},1]$, followed by a normalisation of $\tilde{g}$. It holds that
$\tilde{g}(\tau)=C\,g\big{(}\tfrac{a}{b}+\tfrac{1-a/b}{b-a}(\tau-a)\big{)}$
with a normalising constant $C$. The degree of a polynomial $g$ is invariant in this transformation.
The equivalent system is arranged for the scaled DDE (13). We obtain
$\begin{split}\tilde{y}^{\prime}(\tilde{t})&=b\,f\Big{(}\tilde{y}(t),\sum_{i=0}^{n}\tilde{\alpha}_{i}x_{i}(t)\Big{)}\\ x_{0}^{\prime}(t)&=y(t-\tfrac{a}{b})-y(t-1)\\[6.45831pt] x_{i}^{\prime}(t)&=y(t-\tfrac{a}{b})\left(\tfrac{a}{b}\right)^{i}-y(t-1)\quad\mbox{for}\;\;i=1,\ldots,n\,,\end{split}$ (14)
where the coefficients $\tilde{\alpha}_{i}$ result from the weight function $\tilde{g}$.
4 Approximate DDEs from quadrature
An approach to solve an IVP of a DDE with distributed delay consists in replacing the involved integral by a quadrature formula. The DDE (1) changes into
$y^{\prime}(t)=f\Big{(}y(t),\sum_{k=1}^{m}\omega_{k}y(t-\tau_{k})\Big{)}$ (15)
including the (pairwise different) nodes $\{\tau_{1},\tau_{2},\ldots,\tau_{m}\}\subset I$ and the real-valued weights $\{\omega_{1},\omega_{2},\ldots,\omega_{m}\}$ belonging to a quadrature formula. Thus Eq. (15) represents a DDE with $m$ discrete delays. In [11], the integral was discretised by the (composite) Simpson rule. Initial values for the DDE (1) with distributed delay yield initial conditions for the DDE (15) with discrete delays, since the nodes are located inside the interval $I$.
Now numerical methods for IVPs of DDEs with discrete delays can be applied to (15). An exact solution of a DDE (15) represents an approximation of a solution of the DDE (1) with distributed delay. However, the computation work increases with the number of discrete delays. Hence the number $m$ should be kept small.
Gaussian quadrature can be used for weighted integrals
$J(f)=\int_{I}f(t)\,g(t)\;\mathrm{d}t,$ (16)
where the weight function $g:I\rightarrow\mathbbm{R}_{0}^{+}$ satisfies some assumptions, see [14]. There is an associated Gaussian quadrature scheme for each traditional probability distribution, where the weight function represents a probability density function. Table 1 lists some of these quadrature schemes. Each Gaussian quadrature scheme features an optimality property concerning polynomials $f$ in (16). Given a fixed number $m$ of nodes, the Gaussian quadrature exhibits a maximum possible degree of polynomial exactness $12m-$ 1, i.e.,
$\int_{I}t^{i}g(t)\;\mathrm{d}t=\sum_{k=1}^{m}\omega_{k}\tau_{k}^{i}\qquad\mbox{for}\;\;i=0,1,\ldots,2m-1.$
Thus Gaussian quadrature methods are appropriate for the discretisation of distributed delays in DDEs.
Table 1: Gaussian quadrature rules for weight functions associated to traditional probability distributions.
distribution range $I$ name
uniform distribution $[a,b]$ Gauss-Legendre quadr.
beta distribution $[a,b]$ Gauss-Jacobi quadr.
exponential distribution $[0,\infty)$ Gauss-Laguerre quadr.
gamma distribution $[0,\infty)$ generalised Gauss-Laguerre quadr.
5 Numerical results
Epidemiological models often consist of ODEs or DDEs, whose solutions include populations of susceptibles $S$, infected $I$, and recovered $R$. In [15], an $SIR$ model was investigated including a birth rate and a death rate, where the loss of a temporary immunity is described by a single discrete delay. In [10], the discrete delay was replaced by a distributed delay and numerical solutions were computed using the approach of Section 4. Now we apply this $SIR$ model without birth and death as test example.
Let $y_{1}=S$, $y_{2}=I$, $y_{3}=R$ be the three unknown populations. The system of DDEs with distributed delay reads as
$\begin{split}y_{1}^{\prime}(t)&=-\sigma y_{1}(t)y_{2}(t)+\theta\int_{a}^{b}y_{2}(t-\tau)g(\tau)\;\mathrm{d}\tau\\ y_{2}^{\prime}(t)&=\sigma y_{1}(t)y_{2}(t)-\theta y_{2}(t)\\ y_{3}^{\prime}(t)&=\theta y_{2}(t)-\theta\int_{a}^{b}y_{2}(t-\tau)g(\tau)\;\mathrm{d}\tau\end{split}$ (17)
with infection rate $\sigma>0$ and recovery rate $\theta>0$. It follows that $y_{1}+y_{2}+y_{3}$ is constant in time. We assume the normalisation $y_{1}+y_{2}+y_{3}=1$. The delay appears only in the component $y_{2}$. We apply a beta distribution with probability density function (3) and parameters $p=q=2$. Thus the associated polynomial (2) has degree $n=4$.
We set the other parameters to $\sigma=0.1$ and $\theta=0.05$. The constant initial values $y_{1}(t)=0.99$, $y_{2}(t)=0.01$, $y_{3}(t)=0$ are predetermined for $t\leq 0$. IVPs are solved in the time interval $[0,1000]$. We examine two cases with different ranges of the beta distribution:
  • Case (i): $[a,b]=[30,150]$,
  • Case (ii): $[a,b]=[150,250]$.
To generate the equivalent system with discrete delays for (17), we use functions of type (4)
$x_{i}(t)=\int_{a}^{b}y_{2}(t-\tau)\tau^{i}\;\mathrm{d}\tau$ (18)
for $i=0,1,2,3,4$. The differential equations for (18) are like in (6b), (6c). The initial values follow from (8). In our numerical simulations, the system of DDEs (17) is scaled as specified in Section 3.4. The time intervals are $[0,\frac{20}{3}]$ and $[0,4]$, respectively, in the solution of the scaled systems. Yet the time axis is rescaled to the original time in graphical illustrations.
All numerical computations were executed using the software package Matlab [7]. The built-in routine dde23 [12], which implements an explicit method of Runge-Kutta type, produced the numerical solutions of IVPs of the DDEs with discrete delays. This method features an adaptive step size selection, where we require a relative error tolerance $\varepsilon_{\rm rel}=10^{-6}$ and an absolute error tolerance $\varepsilon_{\rm abs}=10^{-8}$. The integrator performs 610 steps in case (i) and 1058 steps in case (ii). Figure 1 and Figure 2 illustrate the solutions of the IVPs of the equivalent systems (6) in the two cases. In case (i), the solutions tend to a stationary solution. In case (ii), the solutions tend to a periodic oscillation.
(a) populations
(b) additional functions
Figure 1: Numerical solution of DDE system with two discrete delays in case (i).
(a) populations
(b) additional functions
Figure 2: Numerical solution of DDE system with two discrete delays in case (ii).
For comparison, we solve IVPs of the scaled approximate systems of DDEs (15), where Gaussian quadrature is used as described in Section 4. The Gauss-Jacobi quadrature is associated to the beta distribution. In the case of a single node ($m=1$), we obtain a DDE with a single delay, which is identical to the expected value of the probability distribution. This instance represents an alternative model in comparison to the distributed delay. The expected values are $\bar{\tau}=90$ and $\bar{\tau}=200$ in case (i) and case (ii), respectively. The numerical solutions are depicted for the two cases in Figure 3. The solutions qualitatively agree to the situation of distributed delay. Just small quantitative differences are present.
Finally, we determine the differences between the solution of the equivalent system and the solution of the approximate system using different numbers of nodes in the Gauss-Jacobi quadrature. In the time integration, we impose a maximum step size $h_{\max}=10^{-3}$ and $h_{\max}=5\cdot 10^{-4}$ in case (i) and case (ii), respectively. The solutions are evaluated at $1000$ equidistant points in the total time intervals $[0,\frac{20}{3}]$ and $[0,4]$, respectively. Figure 4 shows the maximum differences for components $S$ and $I$, while the differences for $R$ are almost identical to the differences for $S$. We observe that the maximum differences exponentially decay for increasing numbers of nodes. However, the differences stagnate for larger numbers of nodes, because the differences are dominated by the errors of the time integration if the quadrature errors are tiny. These results verify that the equivalent system of the form (6) yields the correct solution of the original DDE (1) in this example.
(a) case (i)
(b) case (ii)
Figure 3: Numerical solution of DDE systems including a single delay.
(a) case (i)
(b) case (ii)
Figure 4: Maximum difference between solutions of equivalent system and solutions of approximate systems using Gaussian quadrature with different numbers of nodes.
6 Conclusions
We considered a DDE with a distributed delay, where the weight function of the distribution is a polynomial. We derived an equivalent system of DDEs including two discrete delays. An exact solution of this system yields an exact solution of the original DDE. We applied a scaling to the DDE with distributed delay, which keeps the magnitude of additional functions small in the equivalent system. However, the equivalent system includes an unstable part. Consequently, errors increase with time in the numerical solution of IVPs. This property implies that IVPs have to be solved with high accuracy to keep the errors sufficiently small. Yet the unstable part is not present in the case of DDEs with a constant weight function. We compared the equivalent system to a DDE with multiple discrete delays, where the integral of the distribution is discretised by Gaussian quadrature. An exact solution of the alternative DDE represents an approximation of an exact solution of the original DDE. We solved IVPs of the DDEs with discrete delays by a numerical integrator. The results of these computations confirm that the equivalent system reproduces solutions of the DDE with distributed delay.